OpenStudy (anonymous):

Find the value of each expression. Show your work. 1. ^ 9P^3 2. ^10C^4

OpenStudy (anonymous):

got a calculator?

OpenStudy (anonymous):

yeah I have but I don't know how to do it

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=+9P3+

OpenStudy (anonymous):

\[_9P_3=9\times 8\times 7\]

OpenStudy (anonymous):

just try your best and you can do it :)

OpenStudy (anonymous):

504

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wat is number 2

OpenStudy (anonymous):

try it and show me your work

OpenStudy (anonymous):

I mean how to to do it

OpenStudy (anonymous):

10C4=5040 / 24 and try it

OpenStudy (anonymous):

210

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

yes 10C4 = Combinations any 4 from 10 @satellite73 is this ok

OpenStudy (anonymous):

\[_{10}C_4=\frac{10\times 9\times 8\times 7}{4\times 3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

that is what I did

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

\[10\times 3\times 7\] you do in your head

OpenStudy (anonymous):

210